OpenStudy (anonymous):

What is the 46th term of the arithmetic series 1, 8, 15, … ?

hero (hero):

I don't know why, but I can never remember the formula for this

OpenStudy (wasiqss):

2(1)+(45)7

OpenStudy (anonymous):

this is an arithmetic progression, and the formula for find an n term is : an = a1 + (n-1)d. so, an = a1[first term, so 1] + (n[number of terms, so 46]-1)d[7]. so: an=1+(46-1)*7 -> an=1 + 45*7 -> an = 315.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

correcting, 316, i forgot adding 1.